``Before, solving the problem, I just want to give you the heads up. It's
Systems of Equation: Application. I can't find the specific title to relate
with my story problems. My chapter in the book is Chapter Eight, Solving Systems by Substiution, Additional. Hope it helps!
*The difference of two integers is 20. The larger is 4 less than 3 times the smaller. Find the two integers.*

Answer by nerdybill(7384)

*The difference of two integers is 20. The larger is 4 less than 3 times the smaller. Find the two integers.*
.
Let x = larger of the two integers
and y = smaller of the two integers
.
x-y=20 (equation 1)
x = 3y-4 (equation 2)
.
Substitute the value of x from equation 2 into equation 1 and solve for y:
x-y=20
3y-4-y=20
2y-4=20
2y=24
y = 12
.
Substitute the value of y above into equation 2 and solve for x:
x = 3y-4
x = 3(12)-4
x = 36-4
x = 32
.
Our two integers are 12 and 32